# Physics-Aware Sparse Signal Recovery Through PDE-Governed Measurement Systems

### v1

@@ V-A Signal Recovery Performance of PA-ISTA
We conducted extensive numerical experiments to evaluate the signal recovery performance of PA-ISTA. The simulations were performed using the following parameter setting. For the NLSE parameters, we set the dispersion constant $\beta_{2}$ to $10$ and the nonlinear coefficient $\gamma$ to $2$. The pulse half-width $T_{0}$ was set to 1, resulting in a dispersion length $L_{D}\equiv T_{0}^{2}/|\beta_{2}|=0.1$, and nonlinear length $L_{NL}\equiv 1/\gamma=0.5$. These parameters were chosen to represent typical propagation conditions where both dispersion and nonlinear effects significantly influence the signal evolution. The SSFM solver was configured with a spatial step size ($z$-direction) of $l=0.01$ and a temporal step size ($t$-direction) of $h=0.3$. The temporal grid consisted of $m=256$ points spanning the interval $[-38.4,38.4]$, providing sufficient resolution for accurate waveform propagation simulation. For our experiments, we generated sparse test vectors with the following parameters. Each test vector had length $n=30$ with exactly $k=3$ non-zero elements, representing a sparse signal. The positions of the non-zero elements were randomly distributed across the vector following a uniform distribution. The values of the non-zero elements were complex numbers, each with unit magnitude ($|s_{i}|=1$) but with uniformly randomly chosen phases.
@@
The signal propagation and measurement were simulated over a distance of $L=5L_{D}=0.5$, requiring 50 iterations of the SSFM solver. To generate received signals, we synthesized the received waveform using the SSFM solver with the previously specified parameters. This approach allows us to evaluate the recovery performance under well-controlled conditions while maintaining the essential nonlinear characteristics of optical fiber propagation. We used $q_{i}=-38.4+ih(i=0,1,\ldots,511)$ as sensing positions. The simulation code is implemented using Julia 1.9 and AD mechanism in Zygote.jl.
@@
The deep unfolding training was conducted under the following experimental conditions. We set the signal-to-noise ratio to $SNR\equiv 1/\sigma^{2}=5$ dB. The network was trained for 100 iterations, with PA-ISTA configured to perform 30 iterations per forward pass. The step size parameters $\eta^{(k)}$ were initialized to $0.01$, and the shrinkage parameters $\theta^{(k)}$ were initialized to $0.001$ for $k=0,1,\ldots,29$. These initial parameters were found by an ad hoc manner. For optimization of the learnable parameters, we employed the Adam optimizer [18] with a learning rate of $10^{-4}$ to simultaneously update both $\eta^{(k)}$ and $\theta^{(k)}$. The deep unfolding architecture was implemented following the methodology described in [17].
@@
Figure 2 presents an example of sparse signal recovery using PA-ISTA with parameters optimized through deep unfolding. Panel (c) shows the recovered signal, which closely matches the original input waveform shown in panel (a), demonstrating the effectiveness of our approach. For comparison, panel (d) shows the result of conventional backpropagation-based recovery $BP(\bm{y})$. The comparison clearly illustrates that backpropagation fails to provide robust recovery in the presence of observation noise, whereas PA-ISTA maintains reliable performance under these challenging conditions.
@@
Figure 2: Example of signal recovery. PA-ISTA with DU-optimized parameters were used (Blue:real part, Red: imaginary part). The SNR was set to 5 dB.
@@
Figure 4: Comparisons of MSE performance of PA-ISTA. SNR is 5 dB.
@@
Figure 4 compares the mean squared error (MSE) performance of PA-ISTA under different SNR conditions (5dB). The MSE is defined as ${\sf E}\left[\|\bm{s}-\hat{\bm{s}}\|_{2}^{2}\right]$, where $\bm{s}$ represents the true sparse vector and $\hat{\bm{s}}$ is the PA-ISTA estimate. The expectation is estimated from 100 independent trials. The results demonstrate that PA-ISTA achieves substantially lower MSE values compared to conventional backpropagation methods, whose performance levels are indicated by horizontal lines in the figure. This significant performance gap clearly highlights the advantages of our physics-aware approach over traditional recovery methods. The figure also includes MSE curves for PA-ISTA using fixed initial parameters without DU optimization. These curves exhibit notably slower convergence, as evidenced by their shallower slopes compared to the versions with optimized parameters. This comparison provides compelling evidence for the effectiveness of DU in enhancing reconstruction performance. The optimization of iteration-dependent parameters through DU leads to both faster convergence and better final recovery accuracy.
V-B Estimation of NLSE Parameters
Figure 5 illustrates the convergence behavior of the gradient descent process for estimating the NLSE parameters $\beta_{2}$ and $\gamma$ through the squared error minimization defined in (20). The initial values were set to $\beta_{2}=8$ and $\gamma=5$, with all other simulation parameters matching those of our previous experiments. The estimation was performed under an SNR of $5$ dB. The trajectories shown in Figure 5 demonstrate gradual convergence of the estimated parameters toward their true values as the iterations progress, exemplifying a successful case of parameter estimation. The smooth convergence behavior suggests that when initialized appropriately, the gradient descent process can effectively navigate the optimization landscape to find accurate parameter estimates.
However, it is important to note that the success of this estimation process is sensitive to initialization. Our experiments revealed that estimation attempts occasionally fail when the initial parameter values are too far from their true values. This sensitivity can be attributed to the non-convex nature of the optimization problem, where the gradient descent process may become trapped in undesirable stationary points. These observations underscore the critical importance of appropriate parameter initialization for achieving reliable estimation results. Accurate calibration of the NLSE parameters is crucial for achieving optimal performance in physics-aware sparse signal recovery, as the reconstruction quality heavily depends on the precision of the underlying physical model.
Figure 5: Parameter estimation via squared error minimization.
VI Concluding summary
@@
This paper has presented a physics-aware framework for sparse signal recovery in systems governed by PDEs. By incorporating physical models directly into the recovery process through AD and numerical PDE solvers, our approach demonstrates the potential of physics-aware signal processing. The proposed PA-ISTA algorithm, while illustrated through optical fiber applications, establishes a general methodology for integrating physical constraints into sparse recovery problems.
Several promising directions for future research emerge from this work:
  1. 1.
    Non-Gaussian noises: While our current formulation assumes Gaussian noise and employs a quadratic loss function, real-world optical systems often encounter various types of noise with different statistical characteristics, such as shot noise or impulsive noise. Adapting our physics-aware recovery framework to these scenarios would require reformulating the objective function to incorporate appropriate noise models and developing robust estimation techniques.
  2. 2.
    Extension to other physical systems: The framework developed here could be adapted to other PDE-governed measurement systems, such as heat diffusion processes, acoustic wave propagation, or electromagnetic field reconstruction. For example, in case of electromagnetic field, we can employ Finite-Difference Time-Domain (FDTD) method as a PDE solver. Each application would require careful consideration of appropriate numerical solvers and physical constraints.
  3. 3.
    Computational efficiency: Investigation of more efficient numerical PDE solvers and automatic differentiation mechanisms could further improve the practical applicability of physics-aware recovery methods. This includes exploring adaptive step-size selection in PDE solving processes.
  4. 4.
    Real-world applications: Validation of the proposed framework with experimental data from various physical systems would provide valuable insights into its practical limitations and potential improvements.
The integration of physical models with sparse signal recovery techniques represents a promising direction for signal processing in complex physical systems. As computational capabilities continue to advance, physics-aware approaches may become increasingly practical for a wider range of applications in science and engineering.

### v2

@@ V-A Signal Recovery Performance of PA-ISTA
We conducted extensive numerical experiments to evaluate the signal recovery performance of PA-ISTA. The simulations were performed using the following parameter setting. For the NLSE parameters, we set the dispersion constant $\beta_{2}$ to 10-10 and the nonlinear coefficient $\gamma$ to $2$. The pulse half-width $T_{0}$ was set to 1, resulting in a dispersion length $L_{D}\equiv T_{0}^{2}/|\beta_{2}|=0.1$, and nonlinear length $L_{NL}\equiv 1/\gamma=0.5$. These parameters were chosen to represent typical propagation conditions where both dispersion and nonlinear effects significantly influence the signal evolution. The SSFM solver was configured with a spatial step size ($z$-direction) of $\Delta z=0.01$ and a temporal step size ($t$-direction) of $\Delta t=0.3$. The temporal grid consisted of $N_{t}=256$ points spanning the interval [38.4,38.4][-38.4,38.4], providing sufficient resolution for accurate waveform propagation simulation. For our experiments, we generated sparse test vectors with the following parameters. Each test vector had length $n=30$ with exactly $k=3$ non-zero elements, representing a sparse signal. The positions of the non-zero elements were randomly distributed across the vector following a uniform distribution. The values of the non-zero elements were complex numbers, each with unit magnitude ($|s_{i}|=1$) but with uniformly randomly chosen phases.
@@
The signal propagation and measurement were simulated over a distance of $L=3L_{D}=0.3$. To generate received signals, we synthesized the received waveform using the SSFM solver with the previously specified parameters. This approach allows us to evaluate the recovery performance under well-controlled conditions while maintaining the essential nonlinear characteristics of optical fiber propagation. We used $q_{i}=-38.4+ih,i=0,1,\ldots,255$ as sensing positions. The simulation code is implemented using Julia 1.9 and AD mechanism in Zygote.jl.
The deep unfolding training was conducted under the following experimental conditions. We define the complex AWGN as $n_{i}\sim\mathcal{CN}(0,\sigma^{2})$ and the SNR by
$\mathrm{SNR\,[dB]}\equiv 10\log_{10}\!\left(\frac{1}{\sigma^{2}}\right).$
@@
In the following experiments, we set SNR to $15$ dB. The network was trained for 100 iterations, with PA-ISTA configured to perform 30 iterations per forward pass. The step size parameters $\eta^{(k)}$ were initialized to $0.01$, and the shrinkage parameters $\theta^{(k)}$ were initialized to $0.001$ for $k=0,1,\ldots,29$. These initial parameters were found by an ad hoc manner. For optimization of the learnable parameters, we employed the Adam optimizer [25] with a learning rate of $10^{-4}$ to simultaneously update both $\eta^{(k)}$ and $\theta^{(k)}$. The deep unfolding architecture was implemented following the methodology described in [24].
@@
Figure 5 presents an example of sparse signal recovery using PA-ISTA with parameters optimized through deep unfolding. Panel (c) shows the recovered signal, which closely matches the original input waveform shown in panel (a), demonstrating the effectiveness of our approach. For comparison, panel (d) shows the result of conventional DBP-based recovery. The comparison clearly illustrates that DBP fails to provide robust recovery in the presence of observation noise, whereas PA-ISTA maintains reliable performance under these challenging conditions.
@@
Figure 5: Example of signal recovery. PA-ISTA with DU-optimized parameters were used (Blue:real part, Red: imaginary part). The SNR was set to 15 dB.
@@
Figure 7: Comparisons of MSE performance of PA-ISTA.
@@
Figure 7 compares the mean squared error (MSE) performance of PA-ISTA under different SNR conditions (5 dB). The MSE is defined as ${\sf E}\left[\|\bm{s}-\hat{\bm{s}}\|_{2}^{2}\right]$, where $\bm{s}$ represents the true sparse vector and $\hat{\bm{s}}$ is the PA-ISTA estimate. The expectation is estimated from 100 independent trials. The results demonstrate that PA-ISTA achieves substantially lower MSE values compared to conventional DBP-based methods, whose performance levels are indicated by horizontal lines in the figure. This significant performance gap clearly highlights the advantages of our physics-aware approach over conventional recovery methods. The figure also includes MSE curves for PA-ISTA using fixed initial parameters without DU optimization. These curves exhibit notably slower convergence, as evidenced by their shallower slopes compared to the versions with optimized parameters. This comparison provides compelling evidence for the effectiveness of DU in enhancing reconstruction performance. The optimization of iteration-dependent parameters through DU leads to both faster convergence and better final recovery accuracy.
V-B QPSK Detection with PA-ISTA
In this subsection, we present the numerical results of the QPSK detection problem with PA-ISTA. The basic setup of the QPSK detection problem follows the same configuration as in the sparse signal recovery case. The parameters of the NLSE model were set to $\beta_{2}=-10$ and $\gamma=2$. The pulse half-width $T_{0}$ was set to 1, resulting in a dispersion length $L_{D}\equiv T_{0}^{2}/|\beta_{2}|=0.1$, and nonlinear length $L_{NL}\equiv 1/\gamma=0.5$. The SSFM solver was configured with a spatial step size ($z$-direction) of $\Delta z=0.01$ and a temporal step size ($t$-direction) of $\Delta t=0.3$. The temporal grid consisted of $N_{t}=256$ points spanning the interval [38.4,38.4][-38.4,38.4], providing sufficient resolution for accurate waveform propagation simulation. The signal propagation and measurement were simulated over a distance of $L=5L_{D}=0.5$. The length of a QPSK signal vector is $n=15$. Deep unfolding was used to optimize the parameters of PA-ISTA.
The symbol error rate (SER) performance of PA-ISTA is shown in Figure 8. As a benchmark, we also show the SER performance of backpropagation-based recovery $DBP(\bm{y})$. We can observe that PA-ISTA achieves much better SER performance than $DBP(\bm{y})$ in all SNR conditions. The performance gap between PA-ISTA and $DBP(\bm{y})$ can be explained by the use of prior knowledge of the QPSK signal in the shrinkage function. The results indicate that the PA-ISTA can be a promising approach for the QPSK detection problem.
Figure 8: Symbol error rate (SER) performance of PA-ISTA for QPSK detection.
VI Conclusion
@@
This paper has presented a physics-aware framework for sparse signal recovery with PDE-governed measurement systems. By incorporating physical models directly into the recovery process through AD and numerical PDE solvers, our approach demonstrates the potential of physics-aware signal processing. The core idea is also shown to be effecitive in decoding process for binary linear codes in PDE-governed channels as well [27]. The key innovation lies in the seamless integration of physical models with a gradient descent process through AD, enabling the algorithm to utilize gradient information from the PDE solver effectively. The proposed PA-ISTA algorithm, while illustrated through optical fiber applications, establishes a general methodology for integrating physical constraints into sparse recovery problems with PDE-governed measurement systems.
Our numerical experiments with optical fiber channels demonstrate that PA-ISTA significantly outperforms conventional recovery methods in terms of MSE in sparse signal recovery, and symbol error rate performance in QPSK detection. These results validate the practical viability of our physics-aware approach for real-world applications.
The primary future challenge lies in the computational complexity introduced by the double-loop structure of our method. Although using coarser grids can partially address this issue, balancing computational efficiency with solution accuracy remains a crucial challenge for practical implementations. Additionally, extending our framework to handle time-varying physical systems and developing theoretical performance guarantees remain important open problems. As computational capabilities continue to advance, physics-aware approaches may become increasingly practical for a wider range of applications in sensing and communication systems.
Acknowledgment
This work is supported by JST CRONOS (JPMJCS25N5).